OpenStudy (anonymous):

tell wether the expression is equivalent 7x and 7x * 1

OpenStudy (lgbasallote):

do you know the value of 2*1 ?

OpenStudy (anonymous):

1* any value...is same... i mean 1*1 = 100 1*2 = 2 1*100 =100

OpenStudy (anonymous):

so...with that idea 7x *1 = 7x hope this helps..

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

Any number times 1 is always the same number

OpenStudy (anonymous):

For example: 5*1=5

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes